Please help me to solve : Jack usually mows the lawn in 3 hours. Marilyn can mow the same yard in 6 hours. How much time would it take for them to mow the lawn together?
Answer by Edwin McCravy(20086)

Please help me to solve : Jack usually mows the lawn in 3 hours. Marilyn can
mow the same yard in 6 hours. How much time would it take for them to mow the
lawn together?
Make this chart:                 |  Number of   |         | rate in
                                 | lawns mowed  |  hours  | lawns/hr.
---------------------------------------------------------------------
Jack working alone               |              |         | 
Marilyn working alone            |              |         |
Jack and Marilyn working together|              |         |

We want to know the number of hours it will take Jack and Marilyn if they work
together.  Let that be x hours to mow 1 lawn so we fill in x for the hours they
will take and 1 for the number of laws they will mow working together. 



                                 |  Number of   |  no. of | rate in
                                 | lawns mowed  |  hours  | lawns/hr.
---------------------------------------------------------------------
Jack working alone               |              |         | 
Marilyn working alone            |              |         |
Jack and Marilyn working together|      1       |    x    |

Jack mows 1 lawn in 3 hours, so we fill in 1 for the number of lawns
he mows and 3 for the number of hours it takes him:

Similarly, Marilyn mows 1 lawn in 6 hours, so we fill in 1 for the 
number of lawns she mows and 6 for the number of hours it takes her:
  
                                 |  Number of   |  no. of | rate in
                                 | lawns mowed  |  hours  | lawns/hr.
---------------------------------------------------------------------
Jack working alone               |      1       |    3    |   
Marilyn working alone            |      1       |    6    |   
Jack and Marilyn working together|      1       |    x    |   

Next we find the rates in lawns/hr by dividing lawns mowed by hours:


                                 |  Number of   |  no. of | rate in
                                 | lawns mowed  |  hours  | lawns/hr.
---------------------------------------------------------------------
Jack working alone               |      1       |    3    |   1/3
Marilyn working alone            |      1       |    6    |   1/6 
Jack and Marilyn working together|      1       |    x    |   1/x

   The equation comes from:


          Jack's rate + Marilyn's rate = their combined rate

                             1/3 + 1/6 = 1/x

Multiply through by the LCD which is 6x

                                2x + x = 6
                                    3x = 6
                                     x = 2

So it would take them 2 hours working together.   
     
Edwin